OpenStudy (anonymous):

7x-14 over 7x simplify

OpenStudy (anonymous):

take 7 common 7(x-2) / 7x 7 cancels out (x-2)/x

OpenStudy (anonymous):

ty soooo much!!!!!!!

OpenStudy (anonymous):

$$ \frac{7x-14}{7x} = \frac{7x}{7x} - \frac{14}{7x} = 1-\frac{2}{x} $$

OpenStudy (anonymous):

thanks mrx but i need help on a different ? now